OpenStudy (anonymous):

y/m+3=n, solve for y

OpenStudy (anonymous):

to do this first subtract 3 from both sides

OpenStudy (anonymous):

what do you get

OpenStudy (anonymous):

do I subtract 3 from itself and n or itself and y/m?

OpenStudy (anonymous):

from it self

OpenStudy (anonymous):

itself mac auto correct

OpenStudy (anonymous):

you should get y/m=n-3

OpenStudy (anonymous):

thats what I got

OpenStudy (anonymous):

next would I divide m from y/m and n-3

OpenStudy (anonymous):

would the answer be y=n-3/m

OpenStudy (anonymous):

no multiply m

OpenStudy (anonymous):

you should get y=n-3*m.

OpenStudy (anonymous):

i meant multiply

OpenStudy (anonymous):

does * mean y= n-3 over m?

OpenStudy (anonymous):